A rectangular counter top is 16 inches longer than it is wide. If the area of the counter top is 420 square inches, find the length and and the width.
Answer by mananth(16946)

let width be x
length = x+16 inches
.
Area of rectangle A = L*W
420 = x(x+16)
420 = x^2 16x
x^2+16x -420 =0
x^2+30x-14x-420=0
x(x+30)-14(x+30)=0
(x+30)(x-14)=0
x= 14 inches the width.
The length = x+16
= 14+16=30 inches.